# T-FIX: Text-Based Explanations with Features Interpretable to eXperts

### v1

@@ D.2 Supernova
D.3 Politeness
Task.
Understanding how linguistic styles, like politeness, vary across cultures is necessary for building better communication, translation, and conversation-focused systems. holmes2012politeness; havaldar2023multilingual. Today’s LLMs exhibit large amounts of cultural bias havaldar2024building, and understanding nuances in cultural differences can help encourage cultural adaptation in models. We use the holistic politeness dataset from havaldar2025comparingstyleslanguagescrosscultural, which consists of conversational utterances between editors from Wikipedia talk pages, annotated by native speakers from four distinct cultures.
Data Selection & Preprocessing.
We sample 100 examples from the data, balanced equally across classes (rude, slightly rude, neutral, slightly polite, polite) and languages (English, Spanish, Japanese, Chinese).
@@ D.3 Politeness
Explanation Prompt.
We show the prompt in Figure˜A8. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
@@
D.4 Emotion
Task.
Understanding and classifying emotion is important for tasks like therapy, mental health diagnoses, etc. denzin1984understanding. Emotion is often expressed implicitly, and understanding such cues can also aid in building LLM systems that handle implied language understanding well havaldar2025entailed. We use the GoEmotions dataset from demszky2020goemotions, consisting of Reddit comments that have been human-annotated for one of 27 emotions (or neutral, if no emotion is present).
Data Selection & Preprocessing.
We sample 100 examples from the data, balanced equally across 28 emotion classes, including neutral. We additionally ensure the comment is over 20 characters, to remove noisy data points and ensure each comment contains enough information for the LLM to make an accurate classification.
@@ D.4 Emotion
Explanation Prompt.
We show the prompt in Figure˜A9. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
@@
D.5 Laparoscopic Cholecystectomy Surgery.
Task.
The task is to identify the safe and unsafe regions for incision. We used the open-source subset of data from (madani2022artificial), which consists of surgeon-annotated images taken from video frames from the M2CAI16 workflow challenge (stauder2016tum) and Cholec80 (twinanda2016endonet) datasets. This consists of 1015 surgeon-annotated images.
Data Selection & Preprocessing.
We selected the first 100 items from the test set where the safe and unsafe regions were of nontrivial area. Each item has three components: an image of dimensions 640 pixels wide by 360 pixels high, a binary mask of the safe regions of the same dimensions, and a binary mask of the unsafe regions of the same dimensions.
@@
Explanation Prompt.
We show the prompt in Figure˜A10. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
@@
    Calot’s triangle cleared - Hepatocystic triangle must be fully cleared of fat/fibrosis so that its boundaries are unmistakable.
@@
    Cystic plate exposed - The lower third of the gallbladder must be dissected off the liver to reveal the shiny cystic plate and ensure the correct dissection plane.
@@
    Only two structures visible - Only the cystic duct and cystic artery should be seen entering the gallbladder before any clipping or cutting.
@@
    Above the R4U line - Dissection must remain cephalad to an imaginary line from Rouviere’s sulcus to liver segment IV to avoid the common bile duct.
@@
    Safe distance from common bile duct - There should be sufficient distance between the common bile duct and the gallbladder wall to ensure safe dissection.
@@
    Infundibulum start point - Dissection should begin at the gallbladder infundibulum-cystic duct junction to stay in safe tissue planes.
@@
    Subserosal plane stay - When separating the gallbladder from the liver, stay in the avascular subserosal cleavage plane under the serosal fat layer.
@@
    Cystic lymph node guide - Identify the cystic lymph node and clip the artery on the gallbladder side of the node to avoid injuring the hepatic artery.
@@
    No division without ID - Never divide any duct or vessel until it is unequivocally identified as the cystic structure entering the gallbladder.
  10. 10.
    Inflammation bailout - If dense scarring or distorted anatomy obscures Calot’s triangle, convert to a subtotal "fundus-first" approach rather than blind cutting.
  11. 11.
    Aberrant artery caution - Preserve any large or tortuous artery (e.g., a Moynihan’s hump) that might be mistaken for the cystic artery.
@@
D.6 Cardiac Arrest
Task.
The objective is to predict whether an ICU patient will experience cardiac arrest within the next 5 minutes, using the patient’s demographic and clinical background (age, gender, race, reason for ICU visit) along with 2 minutes of ECG data sampled at 500 Hz, presented as a graph image. This framing aligns with cardiology literature, which suggests that short ECG windows (30 seconds to a few minutes) are sufficient for reliable prediction nussinovitch2011reliability. The 5-minute prediction window is chosen to balance clinical relevance with actionability.
Data Selection & Preprocessing.
We use ECG and visit data from the open-source Multimodal Clinical Monitoring in the Emergency Department (MC-MED) Dataset kansal2025mcmed. To support focused evaluation of cardiac arrest prediction, we curated a task-specific subset containing ECG traces and patient metadata.
@@ D.6 Cardiac Arrest
Prior work shows that vital sign abnormalities are predictive of outcomes Candel2022-bo; chen2023multimodal. We defined an alarm at any timestamp where three or more of the following vital signs were outside normal range within a two-minute window—a condition known clinically as decompensation:
@@
Explanation Prompt.
Figure A11 shows the prompt used to generate explanations for predicting whether an ICU patient will experience cardiac arrest within 5 minutes, based on 2 minutes of ECG data along with age, gender, race, and ICU admission reason. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The ECG is provided as a graph image of p-signal values sampled at 500 Hz over a 2-minute window, with labeled axes. While we considered supplying the raw signal as text, the input token limits of current LLMs made this infeasible.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
@@
    Ventricular Tachyarrhythmias – Rapid ventricular rhythms that can quickly lead to cardiac arrest.
@@
    Ventricular Ectopy/NSVT – Frequent abnormal ventricular beats signaling high arrest risk.
@@
    Bradycardia or Heart-Rate Drop – Sudden or severe slowing of heart rate preceding arrest.
@@
    Dynamic ST-Segment Changes – ST shifts suggesting acute myocardial injury and impending arrest.
@@
    Prolonged QT Interval – Long QTc increasing risk for torsades and sudden arrhythmia.
@@
    Severe Hyperkalemia Signs – ECG changes from high potassium predicting arrest, especially among patients on dialysis / end stage renal disease.
@@
    Advanced Age – Older age strongly correlates with higher arrest likelihood.
@@
    Male Sex – Males have a higher overall risk of cardiac arrest.
@@
    Underlying Cardiac Disease – Preexisting heart disease increases arrest susceptibility.
@@
    Critical Illness (Sepsis/Shock) – Severe infections or shock states elevate arrest risk through systemic instability.
@@
D.7 Sepsis
Task.
The goal is to predict whether an emergency department (ED) patient is at high risk of developing sepsis within 12 hours, using Electronic Health Record (EHR) data collected during the first 2 hours of their visit. Each input is a time series of records containing a timestamp, the name of a physiological measurement or medication, and its value.
Data Selection & Preprocessing.
We used data from the publicly available MC-MED dataset kansal2025mcmed and curated a task-specific subset for sepsis prediction.
@@ D.7 Sepsis
Explanation Prompt.
Figure A12 shows the prompt used to generate LLM explanations for sepsis risk prediction. We substitute [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The prompt includes a description of the EHR input format: each time-series record consists of a timestamp, a measurement or medication name, and its value.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
@@
Appendix E Expert Validation
E.1 Mass Maps
E.2 Supernova
E.3 Politeness
E.4 Emotion
E.5 Cardiac
E.6 Sepsis

### v2

@@ G.2 Supernova
G.3 Politeness
Task. Understanding how linguistic styles, like politeness, vary across cultures is necessary for building better communication, translation, and conversation-focused systems. (Holmes, 2012; Havaldar et al., 2023b). Today’s LLMs exhibit large amounts of cultural bias (Havaldar et al., 2024), and understanding nuances in cultural differences can help encourage cultural adaptation in models. We use the holistic politeness dataset from Havaldar et al. (2023a), which consists of conversational utterances between editors from Wikipedia talk pages, annotated by native speakers from four distinct cultures.
Data Selection & Preprocessing. We sample 100 examples from the data, balanced equally across classes (rude, slightly rude, neutral, slightly polite, polite) and languages (English, Spanish, Japanese, Chinese).
@@ G.3 Politeness
Explanation Prompt. We show the prompt in Figure˜A9. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
@@
G.4 Emotion
Task. Understanding and classifying emotion is important for tasks like therapy, mental health diagnoses, etc. (Denzin, 1984). Emotion is often expressed implicitly, and understanding such cues can also aid in building LLM systems that handle implied language understanding well (Havaldar et al., 2025). We use the GoEmotions dataset from Demszky et al. (2020), consisting of Reddit comments that have been human-annotated for one of 27 emotions (or neutral, if no emotion is present).
Data Selection & Preprocessing. We sample 100 examples from the data, balanced equally across 28 emotion classes, including neutral. We additionally ensure the comment is over 20 characters, to remove noisy data points and ensure each comment contains enough information for the LLM to make an accurate classification.
@@ G.4 Emotion
Explanation Prompt. We show the prompt in Figure˜A10. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
@@
G.5 Laparoscopic Cholecystectomy
Task. The task is to identify the safe and unsafe regions for incision. We used the open-source subset of data from (Madani et al., 2022), which consists of surgeon-annotated images taken from video frames from the M2CAI16 workflow challenge (Stauder et al., 2016) and Cholec80 (Twinanda et al., 2016) datasets. This consists of 1015 surgeon-annotated images.
Data Selection & Preprocessing. We selected the first 100 items from the test set where the safe and unsafe regions were of nontrivial area. Each item has three components: an image of dimensions 640 pixels wide by 360 pixels high, a binary mask of the safe regions of the same dimensions, and a binary mask of the unsafe regions of the same dimensions.
@@ G.5 Laparoscopic Cholecystectomy
Explanation Prompt. We show the prompt in Figure˜A11. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
@@
    Calot’s triangle cleared: Hepatocystic triangle must be fully cleared of fat/fibrosis so that its boundaries are unmistakable.
@@
    Cystic plate exposed: The lower third of the gallbladder must be dissected off the liver to reveal the shiny cystic plate and ensure the correct dissection plane.
@@
    Only two structures visible: Only the cystic duct and cystic artery should be seen entering the gallbladder before any clipping or cutting.
@@
    Above the R4U line: Dissection must remain cephalad to an imaginary line from Rouviere’s sulcus to liver segment IV umbilical fissure to avoid the common bile duct. Dissection should be carried out along the inferior edge of the gallbladder (well above the line of safety).
@@
    Infundibulum start point: Dissection can begin at the gallbladder infundibulum-cystic duct junction to stay in safe tissue planes, or at the lateral or medial edges of the gallbladder above Rouviere’s sulcus or along the cystic plate to get mobility of the gallbladder first.
@@
    Peritoneal plane: When separating the gallbladder from the liver, stay in the avascular peritoneal cleavage plane.
@@
    Cystic lymph node (calot’s node) guide: Identify the cystic lymph node and clip the artery on the gallbladder side of the node to avoid injuring the hepatic artery.
@@
    No division without ID: Never divide any duct or vessel until it is unequivocally identified as the cystic structure entering the gallbladder.
@@
    Inflammation bailout: If dense scarring or distorted anatomy obscures Calot’s triangle, convert to open surgery or a fenestrated subtotal approach rather than blind cutting.
@@
G.6 Cardiac Arrest
Task. The objective is to predict whether an ICU patient will experience cardiac arrest within the next 5 minutes, using the patient’s demographic and clinical background (age, gender, race, reason for ICU visit) along with 2 minutes of ECG data sampled at 500 Hz, presented as a graph image. This framing aligns with cardiology literature, which suggests that short ECG windows (30 seconds to a few minutes) are sufficient for reliable prediction (Nussinovitch et al., 2011). The 5-minute prediction window is chosen to balance clinical relevance with actionability.
Data Selection & Preprocessing. We use ECG and visit data from the open-source Multimodal Clinical Monitoring in the Emergency Department (MC-MED) Dataset (Kansal et al., 2025). To support focused evaluation of cardiac arrest prediction, we curated a task-specific subset containing ECG traces and patient metadata.
@@ G.6 Cardiac Arrest
Prior work shows that vital sign abnormalities are predictive of outcomes (Candel et al., 2022; Chen et al., 2023). We defined an alarm at any timestamp where three or more of the following vital signs were outside normal range within a two-minute window—a condition known clinically as decompensation:
@@
Explanation Prompt. Figure A12 shows the prompt used to generate explanations for predicting whether an ICU patient will experience cardiac arrest within 5 minutes, based on 2 minutes of ECG data along with age, gender, race, and ICU admission reason. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6. The ECG is provided as a graph image of p-signal values sampled at 500 Hz over a 2-minute window, with labeled axes. While we considered supplying the raw signal as text, the input token limits of current LLMs made this infeasible.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
@@
    Ventricular Ectopy / NSVT: Runs of non-sustained ventricular tachycardia (NSVT) or frequent premature ventricular contractions may indicate electrical instability in critically ill patients, particularly those with underlying coronary disease or cardiomyopathy. While NSVT is generally considered a more benign form of ventricular tachyarrhythmia, its presence reflects transient abnormal rhythms originating from the lower chambers and does not necessarily always signal impending cardiac arrest.
@@
    Bradycardia or Heart-Rate Drop: The onset of significant bradycardia or a sudden $\geq 30\%$ decline in heart rate is a well-documented precursor to in-hospital cardiac arrest (often preceding pulseless electrical activity or asystole) and should be treated as an alarm sign.
@@
    QRS Widening (Conduction Delay): New or progressive prolongation of the QRS duration on the ECG is an ominous finding in the ICU, often observed in the minutes before cardiac arrest and associated with higher mortality due to deteriorating ventricular conduction.
@@
    Dynamic ST-Segment Changes: Acute ischemic changes on continuous ECG (notably ST-segment elevation or depression) signal low blood flow in the coronary arteries, indicating myocardial infarction or injury, and may precede imminent ventricular fibrillation or cardiac arrest. As blood flow continues to decrease, the ischemic heart can fibrillate or go into VT/VF (i.e., cardiac arrest). Although ST segment changes are common, the link to cardiac arrest is rare but possible.
@@
    QTc Prolongation: Prolongation of the corrected QT (QTc) interval reflects abnormal ventricular repolarization and increases vulnerability to malignant arrhythmias such as torsades de pointes. QTc prolongation may result from medications, electrolyte disturbances (e.g., hypokalemia or hypomagnesemia), or underlying cardiac disease and is a known precursor to sudden cardiac arrest.
@@
    Severe Hyperkalemia Signs: Electrocardiographic signs of severe hyperkalemia (such as peaked T-waves, loss of P-waves, and a widening QRS complex) herald an impending arrest – as potassium levels rise, the ECG may evolve to a sine-wave pattern and typically culminate in ventricular fibrillation or asystole without immediate intervention. Hyperkalemia is a frequent cause of in-hospital cardiac arrest especially among patients on dialysis / end stage renal disease. Looking for signs of hyperkalemia can be important to understand risk of cardiac arrest, especially in selected populations.
@@
    Advanced Age: Increasing age is a major risk factor for cardiac arrest, with events being uncommon in younger patients and substantially more frequent in older ICU populations. Patients over 65 years are more prone to sudden deterioration due to reduced physiologic reserve and a higher burden of cardiovascular disease.
@@
    Male Sex: Male gender is associated with a higher incidence of cardiac arrest, as most cardiac arrests occur in men (with women’s risk rising post-menopause).
@@
    Underlying Cardiac Disease: The presence of serious cardiac conditions – such as coronary artery disease (especially a recent myocardial infarction) or severe heart failure – greatly elevates short-term cardiac arrest risk by creating an electrically and hemodynamically unstable myocardium.
@@
    Critical Illness (Sepsis/Shock): Severe sepsis or septic shock substantially raises the likelihood of cardiac arrest in the near term by causing hypoxia, hypotension, and metabolic derangements that often lead to pulseless electrical activity or asystole.
@@
G.7 Sepsis
Task. The goal is to predict whether an emergency department (ED) patient is at high risk of developing sepsis within 12 hours, using Electronic Health Record (EHR) data collected during the first 2 hours of their visit. Each input is a time series of records containing a timestamp, the name of a physiological measurement or medication, and its value.
Data Selection & Preprocessing. We used data from the publicly available MC-MED dataset (Kansal et al., 2025) and curated a task-specific subset for sepsis prediction.
@@ G.7 Sepsis
Explanation Prompt. Figure A13 shows the prompt used to generate LLM explanations for sepsis risk prediction. We substitute [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6. The prompt includes a description of the EHR input format: each time-series record consists of a timestamp, a measurement or medication name, and its value.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below: